OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (8, 1) (7, 9) Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (8, 1) (7, 9) @Mathematics

OpenStudy (anonymous):

First you need the slope (9-1)/(7-8) = 8/-1 = -8 Then using the point given 1 = -8(8) + b 1 = -64 + b 65 = b y = -8x + 65

OpenStudy (anonymous):

thanks